The number of 25-centavo coin in a purse is 4 less than the number of 50-centavo coins.If the total amount is P8,how many coins are there in all?
Answer by stanbon(75887)

The number of 25-centavo coin in a purse is 4 less than the number of 50-centavo coins.If the total amount is P8,how many coins are there in all?
-------
Equations:
t = f - 4
25t + 50f = 800
------
Substitute to get:
25(f-4) + 50f = 800
75f - 100 = 800
75f = 900
f = 12 (# of 50-cent coins
---
Solve for "t":
t = f -4 = 8 (# of 25-cent coins
=======================
Cheers,
Stan H.